OpenStudy (anonymous):

The volume V of a sphere is a function of its radius r. The formula below describes this relationship. What is the value of V(3), the volume of a sphere with a radius of 3 units? A. 27 B. 12 C. 4 D. 36

OpenStudy (anonymous):

You didnt include the formula, so I will use the one I know: \[V(r)=\frac{4}{3} \pi r^3\]Which can be solved by substitution: \[V(3)=\frac{4}{3} \pi (3)^3=4 \pi (3)^2=4\pi(9)=36\pi\]I'm not sure why the answer doesn't line up, but my guess would be that in the formula that is (supposed to be) given for this problem. The volume's "units" contain a \[\pi\] so the answer would just be 36.

OpenStudy (anonymous):

okay thank u :)

OpenStudy (anonymous):

Your welcome :D, if you found this helpful a medal would be highly appreciated.